# Multifidelity Simulation-based Inference for Computationally Expensive Simulators

### v2

@@ Abstract
Across many domains of science, stochastic models are an essential tool to understand the mechanisms underlying empirically observed data. Models can be of different levels of detail and accuracy, with models of high-fidelity (i.e., high accuracy) to the phenomena under study being often preferable. However, inferring parameters of high-fidelity models via simulation-based inference is challenging, especially when the simulator is computationally expensive. We introduce MF-NPE, a multifidelity approach to neural posterior estimation that leverages inexpensive low-fidelity simulations to infer parameters of high-fidelity simulators within a limited simulation budget. MF-NPE performs neural posterior estimation with limited high-fidelity resources by virtue of transfer learning, with the ability to prioritize individual observations using active learning. On one statistical task with analytical ground-truth and two real-world tasks, MF-NPE shows comparable performance to current approaches while requiring up to two orders of magnitude fewer high-fidelity simulations. Overall, MF-NPE opens new opportunities to perform efficient Bayesian inference on computationally expensive simulators.
Machine Learning, ICML
@@ 1 Introduction
The challenge of extending sampling-based SBI methods like Approximate Bayesian Computation (ABC) (Tavaré et al., 1997; Pritchard et al., 1999) to problems with large numbers of parameters has driven significant advancements (Cranmer et al., 2020). Modern SBI approaches aim for the estimation of the posterior through amortized neural posterior estimation (NPE) (Greenberg et al., 2019; Lueckmann et al., 2017; Papamakarios & Murray, 2016), neural likelihood estimation (Papamakarios et al., 2019), neural ratio estimation (Hermans et al., 2020) and respective sequential variants. These approaches have leveraged recent progress in neural density estimation to improve the scalability and accuracy of SBI, allowing parameter inference in problems with higher dimensionality than was previously achievable (Ramesh et al., 2021; Gloeckler et al., 2024).
Despite these advancements, SBI methods face computational challenges for scenarios involving expensive simulations or high-dimensional parameter spaces, as state-of-the-art methods often require extensive simulation budgets to achieve reliable posterior estimates (Lueckmann et al., 2021).
@@
In this work, we present MF-NPE, a multifidelity approach that improves the efficiency of amortized neural posterior estimation for expensive simulators. MF-NPE reduces the computational burden of posterior estimation by pre-training a neural density estimator on low-fidelity simulations and refining the inference with a smaller set of high-fidelity simulations. Additionally, we extend our method with active learning, facilitating targeted parameter space exploration to effectively enhance high-fidelity posterior estimates given single observations. We focus on multifidelity cases where both models are simulators and where the low-fidelity model is a simplified version of the high-fidelity model, designed based on domain expertise.
We demonstrate that for one statistical task, with known ground-truth posterior, and two computationally expensive neuroscience simulators, MF-NPE can identify the posterior distributions more efficiently than NPE.
The core contributions of our work are:
  • A multifidelity approach for NPE. We propose a transfer learning-based method to integrate low- and high-fidelity simulators for efficient posterior estimation of parameters of high-fidelity simulators.
  • Sequential multifidelity NPE. We propose two active learning schemes that focus computational resources on the most relevant or informative high-fidelity simulations, further improving computational efficiency.
2 Related work
2.1 Multifidelity methods for inference
Multifidelity has been widely explored in the context of likelihood-based inference (Peherstorfer et al., 2018), from maximum likelihood estimation approaches (Maurais et al., 2023) to Bayesian inference methods (Vo et al., 2019; Catanach et al., 2020).
@@ 2.1 Multifidelity methods for inference
Currently, few studies tackle the scenario where the likelihood is not explicitly available (Warne et al., 2022; Prescott et al., 2024; Prescott & Baker, 2021). For instance, some multifidelity sampling-based methods have been proposed within the framework of ABC (Prescott & Baker, 2020) or sequential Monte Carlo ABC (Prescott & Baker, 2021). However, despite their potential for multifidelity inference, these methods inherit limitations of ABC approaches, particularly in high-dimensional parameter spaces, where neural density estimators offer more scalable alternatives to complex real-world problems (Lueckmann et al., 2021).
Here, we combine multifidelity approaches with neural density estimation, enabling its application to inference problems with large numbers of model parameters.
2.2 Transfer learning and simulators
To facilitate learning in a target domain, transfer learning borrows knowledge from a source domain (Panigrahi et al., 2021). This is often done when the target dataset is smaller than the source dataset (Larsen-Freeman, 2013) and has successfully been applied to a range of machine learning tasks, e.g., in computer vision (Jiang & Learned-Miller, 2017; Hussain et al., 2019).
@@ 2.2 Transfer learning and simulators
In the context of numerical simulators, transfer learning approaches have been used to lower the simulation budget, for instance, in CO2 forecasting (Falola et al., 2023), surrogate modeling (Wang et al., 2024) and model inversion with physics-informed neural networks (Haghighat et al., 2021). To the best of our knowledge, the potential of transfer learning for computationally-efficient simulation-based inference has not been realized yet.
2.3 Model misspecification
Our work has connections to the issue of model misspecification, a fundamental and widely studied problem in SBI (Gao et al., 2023; Ward et al., 2022; Huang et al., 2023; Frazier et al., 2024; Adachi et al., 2010; Schmitt et al., 2024; Cannon et al., 2022). The issue arises when the probability over the true observation does not fall within the family of estimated probability distributions defined by a model.
A misspecified model can be seen as a form of low-fidelity model of the data-generating process (the high-fidelity model) since this will often rely on several simplifying assumptions. With this perspective in mind, ROPE can be considered a multifidelity approach to SBI (Wehenkel et al., 2024): model misspecification is addressed by providing a small set of real-world observations that re-calibrate through optimal transport the posterior estimates of the misspecified model. In contrast with our work, this scheme does not allow the active sampling of high-fidelity simulations.
@@ 3 Methods
MF-NPE is a multifidelity approach to neural posterior estimation for computationally expensive simulators. We present our approach in Section 3.1. In Section 3.2, we discuss the evaluation metrics used to compare MF-NPE with NPE (Greenberg et al., 2019). MF-NPE is summarized in Algorithm 1 and Fig. 1.
@@ 3.1 Multifidelity NPE
Figure 1: Multifidelity Neural Posterior Estimation (MF-NPE). MF-NPE proceeds by dense sampling from the prior distribution, running the low-fidelity simulator (visualized as a 3D sinusoidal), and training a neural density estimator with a negative log-likelihood loss. MF-NPE then retrains the pre-trained network on sparse samples from the same prior distribution and respective high-fidelity simulations (visualized as a falling waterdrop simulator with exponential decay). Given empirical observations $\boldsymbol{x_{o}}$, MF-NPE estimates the posterior distribution given the high-fidelity model. In the sequential case, the parameters for high-fidelity simulations are drawn from iterative refinements of the prior distribution within the support of the current posterior estimate, at some observation $\boldsymbol{x_{o}}$.
We aim to infer the posterior distribution over the parameters $\boldsymbol{\theta}$ of a computationally expensive high-fidelity simulator $p(\boldsymbol{x}|\boldsymbol{\theta})$, with computational cost $c$. We designate the simulator as high-fidelity if the model accurately captures the empirical phenomenon, but incurs high computational cost when generating simulations. We assume that we have access to a low-fidelity simulator $p_{L}(\boldsymbol{x}_{\mathrm{L}}|\boldsymbol{\theta})$, describing a simplification of the phenomenon of interest with cost $c_{L}\ll c$. We assume that both simulators operate over the same domain of observations $\boldsymbol{x}$, and the parameters of the low-fidelity model form at least a subset (and at most the entirety) of the high-fidelity parameters. Our goal is to develop an estimator that leverages low-fidelity simulations to infer the posterior distribution over parameters of the high-fidelity model with limited high-fidelity simulations, without access to a tractable likelihood for either simulator.
@@
where $\boldsymbol{\theta}$ is sampled from the prior distribution and $\boldsymbol{x}$ denotes the respective simulations (i.e., samples from $p(\boldsymbol{x}|\boldsymbol{\theta})$), and $\boldsymbol{\phi}$ are the network parameters. By minimizing $\mathcal{L}$, the neural density estimator approximates the conditional distribution $p(\boldsymbol{\theta}|\boldsymbol{x})$ directly (Papamakarios & Murray, 2016). Given a true observation $\boldsymbol{\boldsymbol{x_{o}}}$, we can then estimate the posterior over parameters $p(\boldsymbol{\theta}|\boldsymbol{x_{o}})$. To ensure $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x_{o}})$ closely approximates the true posterior $p(\boldsymbol{\theta}|\boldsymbol{x_{o}})$, the density estimator must be sufficiently expressive. We use neural spline flows (NSF) (Durkan et al., 2019), expressive normalizing flows that have been shown empirically to be competitive for SBI (Lueckmann et al., 2021).
@@ 3.1.1 Transfer learning
MF-NPE leverages representations learned from low-fidelity simulations to reduce the number of high-fidelity simulations required to approximate a high-fidelity posterior. To that end, MF-NPE performs transfer learning: MF-NPE pre-trains a neural density estimator $q_{\psi}(\boldsymbol{\theta}|\boldsymbol{x}_{\mathrm{L}})$ on low-fidelity simulations, and uses the network parameters $\psi$ as initialization for a high-fidelity density estimator $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$. We argue that by pre-training on low-fidelity simulations, the density estimator learns useful features up front – i.e., the feature spaces of the low- and high-fidelity density estimators overlap –, so fewer high-fidelity simulations suffice to refine the posterior estimates. Indeed, Tahir et al. (2024) show that once networks learn suitable features for a given predictive task, they drastically reduce the sample complexity for related tasks. Since we do not freeze the weights after transfer, the network retains the flexibility to adapt to high-fidelity simulations without losing the benefits of pre-training.
@@
We apply MF-NPE to cases where the low-fidelity model has fewer parameters than the high-fidelity model. In this setting, the parameters that are exclusive to the high-fidelity model are treated as dummy variables in the pre-trained density estimator, although it would be possible to further improve this pre-conditioning by adapting the network architecture (Papamakarios et al., 2021). We note that the pre-conditioning with the dummy variables leads to the pre-trained neural density estimator to effectively estimate the prior distribution over the respective parameters (see Appendix B.3). As shown below, MF-NPE shows competitive performance compared to NPE, even when the dimensionality of the low- and high-fidelity model are substantially different (e.g., Section 4.3, where the low- and high-fidelity models have 12 and 14 parameters, respectively).
3.1.2 Sequential Training
In addition to learning amortized posterior estimates with NPE, our approach naturally extends to sequential training schemes when estimating the non-amortized posterior $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x_{o}})$. Rather than sampling simulation parameters from the prior, sequential methods introduce an active learning scheme that iteratively refines the posterior estimate for a specific observation $\boldsymbol{x_{o}}$.
These methods – known as Sequential Neural Posterior Estimation (Papamakarios & Murray, 2016; Lueckmann et al., 2017) – have shown increased simulation efficiency when compared to NPE (Lueckmann et al., 2021). However, applying these methods with flexible neural density estimators requires a modified loss that suffers from instabilities in training and posterior leakage (Greenberg et al., 2019). Truncated Sequential Neural Posterior Estimation (TSNPE) mitigates these issues by sampling from a truncated prior distribution that covers the support of the posterior. This leads to a simplified loss function and increased training stability, while retaining performance (Deistler et al., 2022).
We apply our multifidelity approach to TSNPE. First, the high-fidelity density estimator is initialized from the learned network parameters of a low-fidelity density estimator. Then, high-fidelity simulations are generated iteratively from a truncated prior, within the support of the current posterior. We refer to this method as MF-TSNPE (complete description of the algorithm in Appendix F.1).
3.1.3 Acquisition function
To further enhance the efficiency of our sequential algorithm, we explore the use of an acquisition function to supplement our round-wise samples from the TNSPE proposal. In particular, we generate simulations for round $i$ with a set of parameters $\boldsymbol{\theta}^{(i)}=\ \{\boldsymbol{\theta}^{(i)}_{\mathrm{prop}}\cup\boldsymbol{\theta}^{(i)}_{\mathrm{active}}\}$ where $\boldsymbol{\theta}^{(i)}_{\mathrm{prop}}$ are samples from the the proposal distribution at round $i$, and $\boldsymbol{\theta}^{(i)}_{\mathrm{active}}$ are the top $\mathcal{B}$ values according to an acquisition function.
Following Järvenpää et al. (2019); Lueckmann et al. (2019), we select an acquisition function that targets the variance of the posterior estimate with respect to the uncertainty in the learned parameters $\phi|\mathcal{D}$.
We realize this as the sample variance across an ensemble of neural density estimators trained independently on the same dataset $\mathcal{D}$ — as done in Lueckmann et al. (2019). Alternatively, one could estimate $\phi|\mathcal{D}$ using other methods of uncertainty quantification such as Monte Carlo dropout as demonstrated in Griesemer et al. (2024).
The addition of an acquisition function raises the possibility of biasing the proposal distribution, causing the density estimate to diverge from the true posterior. In principle, this could be addressed by using atomic proposals (Greenberg et al., 2019), but given that such an approach suffers from posterior leakage, we do not introduce a proposal correction in order to retain the well-behaved loss function in TNSPE. We hypothesized that the benefit of informative samples would outweigh the potential bias, as long as the percentage of samples selected from the acquisition function would be small compared to the proposal samples, and indeed, our empirical results support this. We refer to this algorithm as A-MF-TSNPE (full description in Appendix F.2).
@@ 3.1.3 Acquisition function
  Input: $N$ pairs of $(\boldsymbol{\theta},\boldsymbol{x}_{\text{L}})$; $M$ pairs of ($\boldsymbol{\theta},\boldsymbol{x})$; conditional density estimators $q_{\psi}(\boldsymbol{\theta}|\boldsymbol{x}_{\text{L}})$ and $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$ with respectively learnable parameters $\psi$ and $\phi$; early stopping criterion $S$.
  $\mathcal{L}(\psi)=\frac{1}{N}\sum_{i=1}^{N}-\log q_{\psi}\left(\boldsymbol{\theta}_{i}|\boldsymbol{x}^{\text{L}}_{i}\right)$
@@
3.2 Evaluation metrics
We evaluate the method on observations $x_{o}$ from the high-fidelity simulator, with parameter values drawn from the prior distribution. This ensured a fair evaluation of how much the low-fidelity simulator helps to infer the posterior distribution given the high-fidelity model.
All methods were evaluated for a range of high-fidelity simulation budgets ($50,10^{2},10^{3},10^{4},10^{5}$), on posteriors given the same data set of observations $x_{o}$.

### v3

@@ Abstract
Across many domains of science, stochastic models are an essential tool to understand the mechanisms underlying empirically observed data. Models can be of different levels of detail and accuracy, with models of high-fidelity (i.e., high accuracy) to the phenomena under study being often preferable. However, inferring parameters of high-fidelity models via simulation-based inference is challenging, especially when the simulator is computationally expensive. We introduce MF-(TS)NPE, a multifidelity approach to neural posterior estimation that uses transfer learning to leverage inexpensive low-fidelity simulations to efficiently infer parameters of high-fidelity simulators. MF-(TS)NPE applies the multifidelity scheme to both amortized and non-amortized neural posterior estimation. We further improve simulation efficiency by introducing A-MF-TSNPE, a sequential variant that uses an acquisition function targeting the predictive uncertainty of the density estimator to adaptively select high-fidelity parameters. On established benchmark and neuroscience tasks, our approaches require up to two orders of magnitude fewer high-fidelity simulations than current methods, while showing comparable performance. Overall, our approaches open new opportunities to perform efficient Bayesian inference on computationally expensive simulators.
@@ 1 Introduction
The challenge of extending sampling-based SBI methods like Approximate Bayesian Computation (ABC) (Tavaré et al., 1997; Pritchard et al., 1999) to problems with large numbers of parameters has driven significant advancements (Cranmer et al., 2020). Modern SBI approaches aim for the estimation of the posterior through amortized neural posterior estimation (NPE) (Greenberg et al., 2019; Lueckmann et al., 2017; Papamakarios & Murray, 2016), neural likelihood estimation (Papamakarios et al., 2019), neural ratio estimation (Hermans et al., 2020; Thomas et al., 2022) and respective sequential variants. These approaches have leveraged recent progress in neural density estimation to improve the scalability and accuracy of SBI, allowing parameter inference in problems with higher dimensionality than was previously achievable (Ramesh et al., 2021; Gloeckler et al., 2024). Despite these advancements, SBI methods face computational challenges for scenarios involving expensive simulations or high-dimensional parameter spaces, as state-of-the-art methods often require extensive simulation budgets to achieve reliable posterior estimates (Lueckmann et al., 2021).
@@
In this work, we present MF-(TS)NPE, a multifidelity approach that improves the efficiency of amortized neural posterior estimation for expensive simulators. MF-(TS)NPE reduces the computational burden of posterior estimation by pre-training a neural density estimator on low-fidelity simulations and refining the inference with a smaller set of high-fidelity simulations. Additionally, we present A-MF-TSNPE, an extension of MF-TSNPE with active learning, facilitating targeted parameter space exploration to effectively enhance high-fidelity posterior estimates given single observations. We focus on multifidelity cases where both models are simulators and where the low-fidelity model is a simplified version of the high-fidelity model, designed based on domain expertise. We demonstrate that for four benchmark tasks and two computationally expensive neuroscience simulators, our multifidelity approach can identify the posterior distributions more efficiently than NPE and TSNPE, often reducing the number of required high-fidelity simulations by orders of magnitude.
2 Background
Multifidelity methods for inference
@@ 2 Background
Multifidelity has been widely explored in the context of likelihood-based inference (Peherstorfer et al., 2018), from maximum likelihood estimation approaches (Maurais et al., 2023) to Bayesian inference methods (Vo et al., 2019; Catanach et al., 2020). For cases where the likelihood is not explicitly available, several sampling-based multifidelity methods have been proposed within the framework of ABC (Prescott & Baker, 2020; Warne et al., 2022; Prescott et al., 2024; Prescott & Baker, 2021). However, these methods inherit limitations of ABC approaches, particularly in high-dimensional parameter spaces, where neural density estimators offer more scalable alternatives to complex real-world problems (Lueckmann et al., 2021). Concurrently with our work, Thiele et al. (2025) developed a multifidelity SBI approach based on response distillation, Hikida et al. (2025) adapted multilevel Monte Carlo techniques to SBI, and Saoulis et al. (2025) applied transfer learning to accelerate inference on a cosmological task.
Beyond SBI, multifidelity has been explored in Bayesian optimization, where Gaussian process models integrate data of different fidelities to infer expensive functions (e.g., Song et al., 2019). These approaches focus on learning surrogate likelihood functions rather than posteriors over simulator parameters, but they highlight the broad applicability of the multifidelity concept.
Transfer learning and simulators
@@
To facilitate learning in a target domain, transfer learning borrows knowledge from a source domain (Panigrahi et al., 2021). This is often done when the target dataset is smaller than the source dataset (Larsen-Freeman, 2013) and has successfully been applied to a range of machine learning tasks, e.g., in computer vision (Jiang & Learned-Miller, 2017; Hussain et al., 2019). In the context of numerical simulators, transfer learning approaches have been used to lower the simulation budget, for instance, in CO2 forecasting (Falola et al., 2023), surrogate modeling (Wang et al., 2024) and model inversion with physics-informed neural networks (Haghighat et al., 2021). To the best of our knowledge, the potential of transfer learning for computationally efficient simulation-based inference has not been fully realized yet.
Simulation-efficient SBI
Recent works have similarly sought to reduce the computational burden of SBI for expensive simulators with active learning frameworks to select informative simulation parameters. Lueckmann et al. (2019) focuses on likelihood estimation, while Griesemer et al. (2024) directly estimates the posterior distribution. Dyer et al. (2022) targets inference under extremely low simulation budgets, employing signature-based feature extraction paired with amortized ratio estimation. (Gutmann & Corander, 2016) proposed an ABC method that uses Bayesian optimisation to efficiently model the discrepancy between the simulator output and the empirical data. More recently, Gloeckler et al. (2025) proposes a compositional framework for time series models that exploits their Markovian structure. Schmitt et al. (2024a) and Schmitt et al. (2024b) introduce self-consistency and consistency models to regularize amortized inference and scale SBI. Delaunoy et al. (2024) show that Bayesian neural networks can be used to obtain calibrated posterior approximations under small simulation budgets. These methods share with MF-NPE the goal of maximizing simulation efficiency through feature extraction and active learning, although they remain within a single-fidelity context.
A related approach involves learning surrogate models to approximate high-fidelity simulators (Wiqvist et al., 2021; Dyer et al., 2023). Both methods train neural surrogate simulators alongside posterior or likelihood estimators, avoiding handcrafted low-fidelity models. In contrast, MF-(TS)NPE assumes the availability of a low-fidelity simulator derived from domain expertise, leveraging transfer learning and active learning strategies to selectively refine posterior estimates.
@@ 3 Methods
MF-(TS)NPE is a multifidelity approach to neural posterior estimation for computationally expensive simulators. We present our approach in Sec. 3.1. In Sec. 3.1.4, we discuss the evaluation metrics used to compare our method against NPE (Greenberg et al., 2019), TSNPE (Deistler et al., 2022), and MF-ABC (Prescott & Baker, 2020). MF-(TS)NPE is summarized in Fig. 1, Algorithms 1 and 3.
@@ 3.1 Multifidelity NPE
Figure 1: Multifidelity Neural Posterior Estimation proceeds by dense sampling from the prior distribution, running the low-fidelity simulator (e.g., a two-compartment neuron model (Hodgkin & Huxley, 1952)), and training a neural density estimator with a negative log-likelihood loss. MF-NPE then retrains the pre-trained network on sparse samples from the same prior distribution and respective high-fidelity simulations (e.g., a multicompartmental neuron model (Rall, 1995)). Given empirical observations $\boldsymbol{x_{o}}$, MF-NPE estimates the posterior distribution given the high-fidelity model. In the sequential case, the parameters for high-fidelity simulations are drawn from iterative refinements of the prior distribution within the support of the current posterior estimate, at some observation $\boldsymbol{x_{o}}$.
We aim to infer the posterior distribution over the parameters $\boldsymbol{\theta}$ of a computationally expensive high-fidelity simulator $p(\boldsymbol{x}|\boldsymbol{\theta})$, with computational cost of a single simulation $c$. We designate the simulator as high-fidelity if the model accurately captures the empirical phenomenon, but incurs high computational cost when generating simulations. We assume that we have access to a low-fidelity simulator $p_{L}(\boldsymbol{x}_{\mathrm{L}}|\boldsymbol{\theta})$, describing a simplification of the phenomenon of interest with cost $c_{L}\ll c$. We assume that both simulators operate over the same domain of observations $\boldsymbol{x}$, and the parameters of the low-fidelity model form at least a subset (and at most the entirety) of the high-fidelity parameters. Our goal is to develop an estimator that leverages low-fidelity simulations to infer the posterior distribution over parameters of the high-fidelity model with limited high-fidelity simulations, without access to a tractable likelihood for either simulator.
@@
where $\boldsymbol{\theta}$ is sampled from the prior distribution, $\boldsymbol{x}$ denotes the respective simulations (i.e., samples from $p(\boldsymbol{x}|\boldsymbol{\theta})$), and $\boldsymbol{\phi}$ are the network parameters. By minimizing $\mathcal{L}(\cdot)$, the neural density estimator approximates the conditional distribution $p(\boldsymbol{\theta}|\boldsymbol{x})$ directly (Papamakarios & Murray, 2016) (proof of convergence in Appendix C). Given an empirical observation $\boldsymbol{x_{o}}$, we can then estimate the posterior over parameters $p(\boldsymbol{\theta}|\boldsymbol{x_{o}})$. To ensure $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x_{o}})$ closely approximates the true posterior $p(\boldsymbol{\theta}|\boldsymbol{x_{o}})$, the density estimator must be sufficiently expressive. We use neural spline flows (NSFs) (Durkan et al., 2019), expressive normalizing flows that have been shown empirically to be competitive for SBI (Lueckmann et al., 2021). To avoid overfitting when training NSFs, we use the same validation-based early stopping criterion $S$ as in the SBI package (Boelts et al., 2024) (details in Appendix.D.1).
@@ 3.1.1 Transfer learning
MF-NPE leverages representations learned from low-fidelity simulations to reduce the number of high-fidelity simulations required to approximate a high-fidelity posterior. To that end, MF-NPE adopts a fine-tuning strategy of transfer learning: Let $\psi$ be the parameters of the low-fidelity neural density estimator $q_{\psi}(\boldsymbol{\theta}|\boldsymbol{x}_{\mathrm{L}})$ and let $\phi$ be the parameters of the high-fidelity density estimator $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$. MF-NPE minimizes the loss $\mathcal{L(\phi)}=\mathbb{E}_{\theta\sim p(\boldsymbol{\theta})}\mathbb{E}_{x\sim p(\boldsymbol{x}|\boldsymbol{\theta})}\left[-\log q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})\right]$ on the high-fidelity task, where the parameters $\phi$ are initialized on the pretrained low-fidelity network parameters $\psi$. We argue that by pre-training on low-fidelity simulations, the density estimator learns useful features up front (i.e., the feature spaces of the low- and high-fidelity density estimators overlap), so fewer high-fidelity simulations suffice to refine the posterior estimates. Indeed, Tahir et al. (2024) shows that once networks learn suitable features for a given predictive task, they drastically reduce the sample complexity for related tasks. Other strategies to pretraining are discussed in Appendix H.3.
@@
MF-NPE can naturally accommodate more than two fidelity levels (Appendix L), does not require more hyperparameter tuning than NPE (Appendix D.1), and is applicable in situations where the low-fidelity model has fewer parameters than the high-fidelity model. In this setting, the parameters that are exclusive to the high-fidelity model are treated as dummy variables in the pre-trained density estimator. The pre-conditioning with these variables leads to the pre-trained neural density estimator to effectively estimate the prior distribution over the respective parameters (OU3 and OU4 in Appendix J.1). As shown below, our method is compatible with both embedding networks and hand-crafted summary statistics of the observations. Overall, MF-NPE shows competitive performance compared to NPE, even when the dimensionality of the low- and high-fidelity models is substantially different (e.g., Sec. 4.3, where the low- and high-fidelity models have 12 and 24 parameters, respectively).
@@
Input: NN pairs of $(\boldsymbol{\theta},\boldsymbol{x}_{\text{L}})$; $M$ pairs of ($\boldsymbol{\theta},\boldsymbol{x})$; conditional density estimators $q_{\psi}(\boldsymbol{\theta}|\boldsymbol{x}_{\text{L}})$ and $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$ with respectively learnable parameters $\psi$ and $\phi$; early stopping criterion $S$.
$\mathcal{L}(\psi)=\frac{1}{N}\sum_{i=1}^{N}-\log q_{\psi}\left(\boldsymbol{\theta}_{i}|\boldsymbol{x}^{\text{L}}_{i}\right)$ .
@@
3.1.2 Sequential training
In addition to learning amortized posterior estimates with NPE, our approach naturally extends to sequential training schemes when estimating the non-amortized posterior $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x_{o}})$. Rather than sampling model parameters from the prior, sequential methods introduce an active learning scheme that iteratively refines the posterior estimate for a specific observation $\boldsymbol{x_{o}}$. These methods – known as Sequential Neural Posterior Estimation (Papamakarios & Murray, 2016; Lueckmann et al., 2017) – have shown increased simulation efficiency when compared to NPE (Lueckmann et al., 2021). However, applying these methods with flexible neural density estimators requires a modified loss that suffers from instabilities in training and posterior leakage (Greenberg et al., 2019). Truncated Sequential Neural Posterior Estimation (TSNPE) mitigates these issues by sampling from a truncated prior distribution that covers the support of the posterior. This leads to a simplified loss function and increased training stability, while retaining performance (Deistler et al., 2022).
We apply our multifidelity approach to TSNPE. First, the high-fidelity density estimator is initialized from the learned network parameters of a low-fidelity density estimator. Then, high-fidelity simulations are generated iteratively from a truncated prior, within the support of the current posterior. We refer to this method as MF-TSNPE (complete description of the algorithm in Appendix M.1).
3.1.3 Acquisition function
To further enhance the efficiency of our sequential algorithm, we explore the use of acquisition functions to supplement our round-wise samples from the TNSPE proposal: we generate simulations for round $i$ with a set of parameters $\boldsymbol{\theta}^{(i)}=\ \{\boldsymbol{\theta}^{(i)}_{\mathrm{prop}}\cup\boldsymbol{\theta}^{(i)}_{\mathrm{active}}\}$ where $\boldsymbol{\theta}^{(i)}_{\mathrm{prop}}$ are samples from the proposal distribution at round $i$, and $\boldsymbol{\theta}^{(i)}_{\mathrm{active}}$ are the top $\mathcal{B}$ values according to an acquisition function. We refer to this algorithm as A-MF-TSNPE (full description in Appendix M.2). Following Järvenpää et al. (2019); Lueckmann et al. (2019), we select an acquisition function that targets the variance of the posterior estimate with respect to the uncertainty in the learned parameters $\phi|\mathcal{D}$.
We realize this as the sample variance across an ensemble of neural density estimators trained independently on the same dataset $\mathcal{D}$, as done in Lueckmann et al. (2019). Alternatively, one could estimate $\phi|\mathcal{D}$ using other methods of uncertainty quantification, such as Monte Carlo dropout (Griesemer et al., 2024). For details on the proposal design of A-MF-TSNPE, see Appendix M.2.
3.1.4 Evaluation metrics
We evaluate the method on observations $\boldsymbol{x_{o}}$ from the high-fidelity simulator, with parameter values drawn from the prior distribution. This ensured a fair evaluation of how much the low-fidelity simulator helps to infer the posterior distribution given the high-fidelity model. All methods were evaluated for a range of high-fidelity simulation budgets ($50,10^{2},10^{3},10^{4},10^{5}$), on posteriors given the same data set of observations $\boldsymbol{x_{o}}$.